(i) Use the inverse D-operator method to solve (i) (D2 – 3D + 5)y = €22. (ii) (D2 + 3)y = sin —2x.
Answer #1

Auxiliary equation m² +p, ma P2 =0 XN 12 If root mazm, then on cfaçem tem mia mza If root mam, & then cfa ot If root afiß the(1) Solution: the given ODE is ²- 3D+5y = 2 To find C.F. to the auxiliary exquation is m²-3m45=0 ..m=t3)2 Vert - 4x145 3219-2The general solution is y=cof. & P.I. g=es (acast Tix &G Sin iTix) + Where a,e constant. are arbitrary(17) Solution: The giver ODE is (1+3) y = Sin (-22) To find cof The auxiliary equation is m²+3=0 m²=-3 m=253 m=oIBi C.F. = eiThe general solution is y = C.F. of P. I > y = ex (e cosbx tQ Sin 13x) - Sin (-28) y = Ci Cos 3x + Sin 3x - Sin2x). Where ca